Circular No. 6205

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

X-RAY NOVA 1994 IN SCORPIUS B. A. Harmon, W. S. Paciesas, and G. J. Fishman, Marshall Space

Flight Center, NASA, report for the Compton Observatory BATSE Team: “The radio-jet transient GRO J1655–40, which has been outburst since late July (IAUC 6196, 6201), reached a peak intensity in the band 20–100 keV of 650 ± 30 mCrab on Aug. 1, decreasing to 80 ± 20 mCrab as of Aug. 9.”

H. Inoue, F. Nagase, and Y. Ueda, on behalf of the ASCA team, communicate: “ASCA will observe this source during Aug. 15.45–16.18 UT, to investigate jet activity. Simultaneous optical and radio observations are encouraged.”

PERSEID METEORS 1995 K. Suzuki, Kozakai Senior High School, Aichi, Japan, reports that an

outburst of Perseid meteors was detected by monitoring forward-scattered meteor echoes via the MU radar wave of Kyoto University (46.5 MHz, 1 MW) at Toyokawa Meteor Observatory in Aichi. The maximum hourly rate of echoes reached 300 during Aug. 12.71–12.79 UT, corresponding to a solar longitude of 139^{o}.6 (equinox 2000.0). The hourly rate of long-period (> 5 s) echoes increased to 5–8 times the usual rate.

COMET C/1995 O1 (HALE-BOPP) D. George, Kanata, ON, provides the following total magnitude esti-

mate from an apparent image of the comet on a prediscovery photograph taken by T. Dickinson (observing in the Chiricahua Mountains, AZ; 85-mm f/1.7 lens, Kodak Royal Gold 400 film): May 29.40 UT, 11.7.

SATURN Further to the item on IAUC 6204, T. Dobbins and S. J. O’Meara

report that a photograph taken with the Lick Observatory 0.91-m refractor on Aug. 12 UT, some 13 hr after ring-plane passage and using an occulting disk to mask the planet’s glare, shows ring material (possibly due to ring E) out to 2.9 Saturn radii on the preceding side only. The first white spot noted on IAUC 6204 had lost contrast (and dimmed) by Aug. 13 UT, and it had drifted to longitude 359^{o} (system I).

1995 August 14 Daniel W. E. Green